OpenStudy (anonymous):

reward and medal, PLEASE HELP;asap Need 2 equations to solve each problem I can do the rest 51.) find tw consecutive integers whose sum is -41.. 52.)find two consecutive even integers e sum is -74 53.)find two consecutive even integers whose sum is 190. 54.) in two consecutive odd integers whose sum is 112.

OpenStudy (mertsj):

51. x+(x+1)=-41 x+(x-1)=-41 52. x+(x+2)=-74 x+(x-2)=-74 53. x+(x+2)=190 x+(x-2)=190 54. x+(x+2)=112 x+(x-2)=112

OpenStudy (anonymous):

the teacher said let 1st= and let 2nd= and something about x alone always equals first I think, not sur... @Mertsj

OpenStudy (anonymous):

so idk if how u did it is what she wants, or if im just misinterpreting it???? please reply

OpenStudy (mertsj):

I don't know either. What does she mean by the first one....the biggest one or what?

OpenStudy (mertsj):

Didn't she give an example?

OpenStudy (anonymous):

idk I think one of your equations is right for everyone but I thin the other part to everyone is suppose to be different, I just don't know how?????????

OpenStudy (mertsj):

51. Let x = first integer Then x+1 = next consecutive integer OR Let x = first integer Then x-1 = integer before the first integer

OpenStudy (mertsj):

Both are ways to represent consecutive integers.

OpenStudy (anonymous):

would I use the add or subtract method u gave please walk me thru 51????????

OpenStudy (mertsj):

\[x+(x+1)=-41\] \[2x+1=-41\] \[2x=-42\] \[x=-21\] \[x+1=-20\]

OpenStudy (mertsj):

The two integers are -20 and -21

OpenStudy (anonymous):

u used x+1 and x-1??

OpenStudy (mertsj):

\[x+(x-1)=-41\] \[2x-1=-41\] \[2x=-40\] \[x=-20\] \[x-1=-20-1=-21\]

OpenStudy (mertsj):

The two integers are -20 and -21

OpenStudy (anonymous):

. what two equations did u use? ur suppose to use two equations.. what two did u use solve problem??????

OpenStudy (mertsj):

For the first problem, I used this equation: x+(x+1)=−41

OpenStudy (mertsj):

For the second problem, I used this equation: x+(x−1)=−41

OpenStudy (anonymous):

do I use those two exzact equations for the rest of problems? just w/a different ending number?????? @Mertsj

OpenStudy (mertsj):

52. x+(x+2)=-74 x+(x-2)=-74 53. x+(x+2)=190 x+(x-2)=190 54. x+(x+2)=112 x+(x-2)=112

OpenStudy (mertsj):

Remember when I posted them at the very beginning when you said just give me the equations I can do the rest?

OpenStudy (anonymous):

yes, thanks(: when u walked m thru 51 I unerstoodnow.. thanks so much @mertsj

OpenStudy (mertsj):

yw